OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

AB=(5-4,2-1) =(1,1) perpendicular vector to (1,1) is for example (1,-1) standart form is: ax+by=c where (a,b) is perpendicular vector to the line. In our case (1,-1) so: 1*(x-4)-1*(y-1)=0 x-4-y+1=0 x-y=4-1 x-y=3 this is your equation